OpenStudy (anonymous):

Help, will medal and fan if givin correct answer ! 1 If f(x) = x2 and g(x) = ----- , find g(f(-1)). 2x + 3

OpenStudy (campbell_st):

start by finding f(-1) which requires you to substitute x = -1 into f(x) then when you get a number answer substitute it into g(x)

OpenStudy (solomonzelman):

no needs for medals, all I need is some effort on your behalf:) I will give you a similar example: g(f(-2)) \(\large\color{black}{ f(x)=x^3 \\g(x)=\frac{\LARGE 1}{\LARGE 3x+4} }\) \(\large\color{black}{ g(\color{red}{f(x)})=\frac{\LARGE 1}{\LARGE 2\color{red}{(x^3)}+3} }\) \(\large\color{black}{ g(\color{red}{f(x)})=\frac{\LARGE 1}{\LARGE 2x^3+3} }\) \(\large\color{black}{ g(\color{red}{f(\color{blue}{-2})})=\frac{\LARGE 1}{\LARGE 2(\color{blue}{-2})^3+3} }\) \(\large\color{black}{ g(\color{red}{f(\color{blue}{-2})})=\frac{\LARGE 1}{\LARGE 2(-8)+3} }\) \(\large\color{black}{ g(\color{red}{f(\color{blue}{-2})})=\frac{\LARGE 1}{\LARGE -16+3} }\) \(\large\color{black}{ g(\color{red}{f(\color{blue}{-2})})=\frac{\LARGE 1}{\LARGE -13} }\)

OpenStudy (anonymous):

So my answr is either 1 or -1 ?

OpenStudy (solomonzelman):

no

OpenStudy (solomonzelman):

~ plug in x^2 instead of x, into the g(x). ~ plug in "-1" instead of x. (once you perform the first step) ~ simplify the denominator. what do you get>?

OpenStudy (campbell_st):

well look at it this way f(-1) = (-1)^2 so what value does f(-1) take..?

OpenStudy (anonymous):

i dont konow, im good at math but this stuff confuses me. im nly n 9th grade

OpenStudy (campbell_st):

well all you are being asked to do is square negative 1... so f(-1) = -1 x -1 what do you think the answer is..?

OpenStudy (anonymous):

multiply them ?

OpenStudy (campbell_st):

yes

OpenStudy (campbell_st):

what answer did you get..?

OpenStudy (anonymous):

1

OpenStudy (campbell_st):

great now this value of substituted into g(x) so f(x) = 1 \[g(1) = \frac{1}{2 \times 1 + 3}\] what answer do you get..?